A mason can make a concrete walk in 6 hours. His helper requires 9 hours to do the same job. In how many hours can they make the walk if they work together?
Answer by nyc_function(2741)

Mason = 1/6
Helper = 1/9
Together = 1/x
1/6 + 1/9 = 1/x
LCD = 18x
3x + 2x = 18
5x = 18
x = 18/5
x = 3.6 hours to work together